### v1

# Equivariant Filter for Tightly Coupled LiDAR-Inertial Odometry*

Abstract
Pose estimation is a crucial problem in simultaneous localization and mapping (SLAM). However, developing a robust and consistent state estimator remains a significant challenge, as the traditional extended Kalman filter (EKF) struggles to handle the model nonlinearity, especially for inertial measurement unit (IMU) and light detection and ranging (LiDAR). To provide a consistent and efficient solution of pose estimation, we propose Eq-LIO, a robust state estimator for tightly coupled LIO systems based on an equivariant filter (EqF). Compared with the invariant Kalman filter based on the $\operatorname{SE}_{2}(3)$ group structure, the EqF uses the symmetry of the semi-direct product group to couple the system state including IMU bias, navigation state and LiDAR extrinsic calibration state, thereby suppressing linearization error and improving the behavior of the estimator in the event of unexpected state changes. The proposed Eq-LIO owns natural consistency and higher robustness, which is theoretically proven with mathematical derivation and experimentally verified through a series of tests on both public and private datasets.
Index Terms:
LiDAR-inertial odometry, Equivariant filter.
I Introduction
In recent years, light detection and ranging (LiDAR) sensors have gained widespread use in simultaneous localization and mapping (SLAM) due to their ability to capture precise depth information. However, LiDAR is vulnerable to distortion caused by rapid movements. In contrast, inertial measurement units (IMUs) can provide motion data at high sampling rates regardless of external environmental conditions. IMU and LiDAR exhibit complementary characteristics, integrating these sensors can significantly enhance the continuity and accuracy of the output [1, 2, 3].
As a result, LiDAR-inertial odometry (LIO) has been widely used in industry. Among various approaches, filter-based methods are well-suited for low-cost platforms with limited computing resources, as they can efficiently handle large volumes of real-time measurements while maintaining lower computational demands. Among filter-based methods, the extended Kalman filter (EKF) achieves high efficiency, albeit with some loss of accuracy. This trade-off makes it a classic data fusion method, as seen in applications like [4] and others. However, researchers have identified that EKF may yield inconsistent estimates, where the computed covariance deviates from the true covariance [5]. To alleviate the inconsistency problem, Barrau and Bonnabel proposed the invariant extended Kalman filter (IEKF) [6]. IEKF ensures the consistency of the bias-free IMU system by analyzing the system’s group affine properties. Later, Shi et al. [7] applied the IEKF for fusing IMU and LiDAR measurements to reduce estimation error, but the IMU bias will destroy the symmetry of the system. Mahony et al. proposed a modern method known as equivariant filtering (EqF) [8, 9], which naturally enhances consistency and robustness with lower computational overhead by incorporating different system symmetries and fixed linearization points. In light of these advantages, our work fully leverages EqF.
In this article, we introduce Eq-LIO, a fast and reliable tightly coupled LIO framework based on the EqF state estimator. Our approach leverages the semi-direct product group to incorporate symmetries, including bias and extrinsic parameters, into the LIO framework. Additionally, we optimize the gravity constraint on the manifold $\mathbb{S}^{2}$. The error dynamics of the EqF ensures that it always performs linearization operations at a fixed origin, which is the key to improving system consistency and reducing linearization errors. Finally, we validate the performance of Eq-LIO in both standard and challenging scenarios. The main contributions of our work are as follows:
  1. 1.
    We introduce Eq-LIO, a tightly coupled LIO system based on the EqF state estimator, with LiDAR and IMU self-calibration capabilities and gravity constraints on $\mathbb{S}^{2}$. To the best of our knowledge, this is the first LIO system to utilize an equivariant filter.
  2. 2.
    We perform extensive benchmark testing across diverse datasets, demonstrating that the proposed framework delivers superior accuracy, robustness, and consistency without increasing computational demands. To support community development, our source code is available online: https://github.com/Eliaul/Eq-LIO.
II Related Works
In this section, we focus on the most related works on LIO and Kalman filtering.
II-A LiDAR-Inertial Odometry
LIO methods can be broadly categorized into two approaches: loosely coupled methods and tightly coupled methods. Loosely coupled methods typically process measurements from the IMU and LiDAR independently before fusing the results. For instance, the LOAM algorithm proposed in [10] utilized the pose estimated from IMU data as an initial guess for LiDAR scan registration. In [11], an adaptive EKF was employed to fuse IMU measurements with poses estimated by LiDAR SLAM, enhancing the accuracy of pose estimation. The direct LiDAR odometry method introduced in [12] enabled real-time, high-speed, and high-precision processing of dense point clouds. While loosely coupled methods offer high computational efficiency, their limited ability to effectively fuse sensor data may lead to unreliable results in the later stages of the fusion process.
The tightly coupled methods directly fuse LiDAR point cloud data with IMU measurements to obtain an optimal state estimation, which can be divided into two primary categories: optimization-based methods and filter-based methods. For optimization-based methods, Ye et al.[1] introduced the LIOM framework, which applied sliding window optimization to LIO. Similarly, LIO-SAM [2] employed smoothing and mapping techniques to achieve sensor fusion and global optimization. In filter-based methods, LINS [3] utilized a robo-centric iterative Kalman filter for tightly coupled pose optimization. FAST-LIO [13] utilized a world-centric iterative Kalman filtering for state estimation, incorporating a novel Kalman gain formula to enhance computational efficiency. FAST-LIO2 [14] further improved system efficiency and robustness by using direct point cloud registration and an ikd-tree structure.
II-B Kalman Filtering for Navigation Applications
Since the introduction of the classic extended Kalman filter (EKF), it has become a widely used tool in filter-based SLAM algorithms. However, the traditional EKF-based LIO algorithm is susceptible to overconfidence issues due to spurious information gain in unobservable directions. To address the inconsistency issues inherent in EKF, Barrau and Bonnabel proposed the invariant extended Kalman filter (IEKF) [6], demonstrating that a consistent filtering algorithm can be developed on group affine systems. This method assumes that the IMU is bias-free, which could not be ignored in practical applications. Despite this limitation, many researchers [15], [16] continue to use this imperfect IEKF framework in the design of SLAM algorithms. They defined the IMU bias on $\mathbb{R}^{3}$, which destroys the group affine property of the system. Recently, SuIn-LIO [17] was proposed, combining an IEKF with an efficient surfel-based map to achieve high-precision LIO. Shi et al.[7] applied the IEKF to the robo-centric and world-centric based LIO and demonstrated that the world-centric method achieves a higher level of accuracy.
Unlike IEKF, which parameterizes the state using Lie groups, Mahony et al. developed an equivariant filtering (EqF) framework [8], [18] that generalizes the state space from Lie groups to homogeneous manifolds. Additionally, EqF does not require constant changes to the linearization point, which naturally ensures system consistency. Subsequently, Fornasier et al. applied EqF to an inertial navigation system [19], which accounts for bias, and demonstrated that various filters, including the EKF, can be incorporated within the EqF framework [20]. Recently, Fornasier et al. [21] introduced an equivariant filter design for vision-assisted INS systems. This new design has been demonstrated to outperform state-of-the-art methods in both robustness and transient behavior.
III Mathematical Preliminaries and Notation
III-A Smooth Manifolds and Lie Theory
Let $\mathcal{M}$ be a smooth manifold, we use $T_{\xi}\mathcal{M}$ to denote the tangent space of $\mathcal{M}$ at point $\xi\in\mathcal{M}$ and $T\mathcal{M}$ to denote the tangent bundle. The notion $\mathfrak{X}(\mathcal{M})$ denotes the set of smooth vector field on $\mathcal{M}$, where each element of $\mathfrak{X}(\mathcal{M})$ assigns a tangent vector in $T_{\xi}\mathcal{M}$ to each point $\xi\in\mathcal{M}$. A Lie group $G$ is a smooth manifold with a group structure, and group multiplication and inversion operations are smooth as mappings between manifolds. For any elements $X,Y\in G$, the group multiplication is denoted by $XY$, the inverse of $X$ is denoted by $X^{-1}$, the identity is denoted by $I$. The tangent space of $G$ at $I$ is called the Lie algebra of $G$, denoted by $\mathfrak{g}$, that is, $\mathfrak{g}=T_{I}G$. The Lie algebra $\mathfrak{g}$ is a vector space with dimension equal to $n=\dim G$. So we have an isomorphism from $\mathfrak{g}$ to $\mathbb{R}^{n}$, which we denote $(\cdot)^{\vee}:\mathfrak{g}\to\mathbb{R}^{n}$, and the inverse of $(\cdot)^{\vee}$ is $(\cdot)^{\wedge}:\mathbb{R}^{n}\to\mathfrak{g}$. In this work, we focus only on the Lie group of matrices, i.e., $G$ is a subgroup of the general linear group $\operatorname{GL}_{n}(\mathbb{R})$.
Fixing a point $X\in G$, there are two important mappings $L_{X}:G\to G$ and $R_{X}:G\to G$, which are called left translation and right translation respectively:
$L_{X}(Y)=XY,\quad R_{X}(Y)=YX.$
Given an element $X\in G$, consider the conjugate action $Y\mapsto XYX^{-1}$ on the Lie group. The differential of this action at the identity $I$ is
$\operatorname{Ad}_{X}:\mathfrak{g}\to\mathfrak{g},\quad\operatorname{Ad}_{X}(\boldsymbol{u})=(dL_{X})\circ(dR_{X^{-1}})(\boldsymbol{u}),$
for every $\boldsymbol{u}\in\mathfrak{g}$. This map $\operatorname{Ad}_{X}$ is called a (big) Adjoint map. Since we have an isomorphism $(\cdot)^{\vee}:\mathfrak{g}\to\mathbb{R}^{n}$, $\operatorname{Ad}_{X}^{\vee}:=(\cdot)^{\vee}\circ\operatorname{Ad}_{X}\circ(\cdot)^{\wedge}$, as a linear map $\mathbb{R}^{n}\to\mathbb{R}^{n}$, can be viewed as a matrix.
For any $u\in\mathfrak{g}$, the (little) adjoint map is defined by
$\operatorname{ad}_{u}:\mathfrak{g}\to\mathfrak{g},\quad\operatorname{ad}_{u}(v)=[u,v],$
where $[u,v]$ is the Lie bracket.
III-B Group Action, Useful Maps and Notation Explanation
Assuming $G$ is a Lie group and $\mathcal{M}$ is a smooth manifold, we consider the Lie group action $\phi:G\times\mathcal{M}\to\mathcal{M}$. Fixing an element $X\in G$, we write $\phi_{X}:\mathcal{M}\to\mathcal{M}$ to represent
$\phi_{X}(\xi):=\phi(X,\xi).$
Fix an element $\xi\in\mathcal{M}$, we write $\phi^{(\xi)}:G\to\mathcal{M}$ to represent
$\phi^{(\xi)}(X):=\phi(X,\xi).$
For all $\boldsymbol{X}=(\boldsymbol{C},\boldsymbol{a},\boldsymbol{b})\in\operatorname{SE}_{2}(3)=\operatorname{SO}(3)\ltimes(\mathbb{R}^{3}\oplus\mathbb{R}^{3})$, define $\Gamma$ to extract the rotation part of $\boldsymbol{X}$:
$\Gamma:\operatorname{SE}_{2}(3)\to\operatorname{SO}(3),\quad\Gamma(\boldsymbol{X})=\boldsymbol{C}.$
In this article, the use of bold and non-bold letters for the same character represents two different meanings. Bold letters are generally used to denote matrices, while non-bold letters are typically used to represent elements in groups or manifolds.
IV System Overview
Figure 1: Overview of Eq-LIO. The equivariant filtering state estimator is employed in the odometry module, with the measurement model constructed through scan-to-map matching. The red box in the figure highlights the main content of this paper. For detailed information, please refer to Section IV and V.
The pipeline of the proposed Eq-LIO is illustrated in Fig. 1. The raw point cloud is first de-skewed using IMU predicted pose [14]. Next, a world-centric equivariant filter state estimator is constructed for the odometry module, utilizing the scan-to-map matching method to establish correspondence between the LiDAR scan and the global map. Finally, the output from the state estimator is used to update the global map.
IV-A System Definition
The IMU coordinate system at the origin is used as the reference coordinate system, which we call the world coordinate system and is represented by $w$. We use $b$ to represent the body coordinate system. The system model of LIO is as follows.
$\dot{\boldsymbol{C}}_{b}^{w}$ $=\boldsymbol{C}_{b}^{w}(\boldsymbol{\omega}-\boldsymbol{b}_{g})^{\wedge},$ (1)
$\dot{\boldsymbol{v}}_{wb}^{w}$ $=\boldsymbol{C}_{b}^{w}(\boldsymbol{a}-\boldsymbol{b}_{a})+\boldsymbol{g}^{w},$
$\dot{\boldsymbol{r}}_{wb}^{w}$ $=\boldsymbol{v}_{wb}^{w},$
$\qquad\dot{\boldsymbol{b}}_{g}$ $=\boldsymbol{0}_{3\times 1},\quad\dot{\boldsymbol{b}}_{a}=\boldsymbol{0}_{3\times 1},$
$\dot{\boldsymbol{C}}^{b}$ $=\boldsymbol{O}_{3\times 3},\quad\dot{\boldsymbol{l}}^{b}=\boldsymbol{0}_{3\times 1},$
where $\boldsymbol{C}_{b}^{w}$, $\boldsymbol{v}_{wb}^{w}$, $\boldsymbol{r}_{wb}^{w}$ represents the attitude, velocity and position of the IMU, $\boldsymbol{b}_{g}$ and $\boldsymbol{b}_{a}$ represent the gyroscope bias and accelerometer bias, respectively, and $\boldsymbol{C}^{b}$ and $\boldsymbol{l}^{b}$ represent the attitude and position of the LiDAR relative to the IMU, $\boldsymbol{g}^{w}$ represents the gravity vector, $\boldsymbol{w}$ and $\boldsymbol{a}$ represent the angular velocity and specific force measured by the IMU.
We introduce a virtual “velocity bias” $\boldsymbol{b}_{\mu}$ and virtual inputs $\boldsymbol{\mu},\boldsymbol{\tau}_{c},\boldsymbol{\tau}_{g},\boldsymbol{\tau}_{a},\boldsymbol{\tau}_{\mu},\boldsymbol{\tau}_{l}$ to exploit the geometric properties of the system, the equation (1) expands to [19]:
$\dot{\boldsymbol{C}}_{b}^{w}$ $=\boldsymbol{C}_{b}^{w}(\boldsymbol{\omega}-\boldsymbol{b}_{g})^{\wedge},$ $\qquad\dot{\boldsymbol{b}}_{g}$ $=\boldsymbol{\tau}_{g},$ (2)
$\dot{\boldsymbol{v}}_{wb}^{w}$ $=\boldsymbol{C}_{b}^{w}(\boldsymbol{a}-\boldsymbol{b}_{a})+\boldsymbol{g}^{w},$ $\dot{\boldsymbol{b}}_{a}$ $=\boldsymbol{\tau}_{a},$
$\dot{\boldsymbol{r}}_{wb}^{w}$ $=\boldsymbol{C}_{b}^{w}(\boldsymbol{\mu}-\boldsymbol{b}_{\mu})+\boldsymbol{v}_{wb}^{w},$ $\dot{\boldsymbol{b}}_{\mu}$ $=\boldsymbol{\tau}_{\mu},$
$\dot{\boldsymbol{C}}^{b}$ $=\boldsymbol{C}^{b}\boldsymbol{\tau}_{c}^{\wedge},$ $\dot{\boldsymbol{l}}^{b}$ $=\boldsymbol{\tau}_{l}.$
The system state is modeled as $\xi=(\boldsymbol{T},\boldsymbol{b}^{\wedge},\boldsymbol{K})\in\mathcal{M}$, with navigation state $\boldsymbol{T}=(\boldsymbol{C}_{b}^{w},\boldsymbol{v}_{wb}^{w},\boldsymbol{r}_{wb}^{w})\in\operatorname{SE}_{2}(3)$, bias $\boldsymbol{b}=(\boldsymbol{b}_{g},\boldsymbol{b}_{a},\boldsymbol{b}_{\mu})\in\mathbb{R}^{9}$ and extrinsic parameters $\boldsymbol{K}=(\boldsymbol{C}^{b},\boldsymbol{l}^{b})\in\operatorname{SE}(3)$ of LiDAR and IMU. Let $\boldsymbol{u}=(\boldsymbol{w}^{\wedge},\boldsymbol{g}^{\wedge},\boldsymbol{\tau}^{\wedge},\boldsymbol{\tau}_{k}^{\wedge})\in(\mathfrak{se}_{2}(3))^{3}\times\mathfrak{se}(3)$ be the system inputs, where $\boldsymbol{g}^{\wedge}=(\boldsymbol{0}^{\wedge},\boldsymbol{g}^{w},\boldsymbol{0})\in\mathfrak{se}_{2}(3)$, $\boldsymbol{\tau}^{\wedge}=(\boldsymbol{\tau}_{g}^{\wedge},\boldsymbol{\tau}_{a},\boldsymbol{\tau}_{u})\in\mathfrak{se}_{2}(3)$, $\boldsymbol{\tau}_{k}^{\wedge}=(\boldsymbol{\tau}_{c}^{\wedge},\boldsymbol{\tau}_{l})\in\mathfrak{se}(3)$ and $\boldsymbol{w}=(\boldsymbol{\omega},\boldsymbol{a},\boldsymbol{\mu})$ be the IMU measurements. In general, the system state evolves on the manifold $\mathcal{M}=\operatorname{SE}_{2}(3)\times\mathfrak{se}_{2}(3)\times\operatorname{SE}(3)$, and the system inputs are in the input space $L=(\mathfrak{se}_{2}(3))^{3}\times\mathfrak{se}(3)$. In this notation, the system model (2) can be rewritten as
$\dot{\boldsymbol{T}}$ $=f_{1}^{0}(\boldsymbol{T})\boldsymbol{T}+\boldsymbol{T}(\boldsymbol{w}^{\wedge}-\boldsymbol{b}^{\wedge})+\boldsymbol{g}^{\wedge}\boldsymbol{T},$ (3)
$\dot{\boldsymbol{b}}$ $=\boldsymbol{\tau},$
$\dot{\boldsymbol{K}}$ $=\boldsymbol{K}\boldsymbol{\tau}_{k}^{\wedge},$
where $f_{1}^{0}:\operatorname{SE}_{2}(3)\to T\operatorname{SE}_{2}(3)$ defined by $\boldsymbol{X}=(\boldsymbol{C},\boldsymbol{v},\boldsymbol{r})\mapsto(\boldsymbol{O},\boldsymbol{0},\boldsymbol{v})\in T_{\boldsymbol{X}}\operatorname{SE}_{2}(3)$ is a vector field on $\operatorname{SE}_{2}(3)$. We can write (3) in a more compact form:
$\dot{\xi}$ $=f^{0}(\xi)+f_{\boldsymbol{u}}(\xi)$ (4)
$=f^{0}(\xi)+\bigl{(}\boldsymbol{T}(\boldsymbol{w}-\boldsymbol{b})^{\wedge}+\boldsymbol{g}^{\wedge}\boldsymbol{T},\boldsymbol{\tau}^{\wedge},\boldsymbol{K}\boldsymbol{\tau}_{k}^{\wedge}\bigr{)},$
where $f^{0}:\mathcal{M}\to T\mathcal{M}$ defined by $\xi\mapsto(f_{1}^{0}(\boldsymbol{T}),\boldsymbol{O},\boldsymbol{O})\in T_{\xi}\mathcal{M}$ is a vector field on $\mathcal{M}$.
IV-B Equivariant Symmetry of the System
The symmetry of the system arises from the key semidirect product group $G=\bigl{(}\operatorname{SE}_{2}(3)\ltimes\mathfrak{se}_{2}(3)\bigr{)}\times\operatorname{SE}(3)$. Let $X=(A,a^{\wedge},B)\in G$, $\gamma=(\gamma_{1}^{\wedge},\gamma_{2}^{\wedge},\gamma_{3}^{\wedge},\gamma_{4}^{\wedge})\in L$.
Proposition IV.1
Define $\phi:G\times\mathcal{M}\to\mathcal{M}$ as
$\phi(X,\xi)=(\boldsymbol{T}A,\operatorname{Ad}_{A^{-1}}(\boldsymbol{b}^{\wedge}-a^{\wedge}),\Gamma(A)^{-1}\boldsymbol{K}B),$ (5)
where $\operatorname{Ad}_{A^{-1}}:\mathfrak{se}_{2}(3)\to\mathfrak{se}_{2}(3)$ represents the isomorphism induced by conjugation $Y\mapsto A^{-1}YA$ on $\operatorname{SE}_{2}(3)$. Then $\phi$ is a transitive and free right group action of $G$ on $\mathcal{M}$ [22].
Proposition IV.2
Define $\psi:G\times L\to L$ as
$\psi(X,\gamma)=\bigl{(}\operatorname{Ad}_{A^{-1}}(\gamma_{1}^{\wedge}-a^{\wedge})+f_{1}^{0}(A^{-1}),\\ \gamma_{2}^{\wedge},\operatorname{Ad}_{A^{-1}}\gamma_{3}^{\wedge},\operatorname{Ad}_{B^{-1}}\gamma_{4}^{\wedge}\bigr{)},$ (6)
then $\psi$ is a right group action of $G$ on $L$.
Theorem IV.3
System (4) is equivariant with group action in IV.1 and IV.2, that is,
$f^{0}(\xi)+f_{\psi_{X}(\gamma)}(\xi)\\ =d\phi_{X}\bigl{(}f^{0}(\phi_{X^{-1}}(\xi))+f_{\gamma}(\phi_{X^{-1}}(\xi))\bigr{)}$ (7)
for every $X\in G,\xi\in\mathcal{M}$ and $\gamma\in L$.
IV-C Lifted System
Equivariant filtering requires a lift $\Lambda:\mathcal{M}\times L\to\mathfrak{g}$ to transfer the system differential equations to the Lie group $G$. The lift $\Lambda$ requires
$d\phi^{(\xi)}\circ\Lambda(\xi,\gamma)=f_{\gamma}(\xi),$
where $\xi\in\mathcal{M}$ and $\gamma\in L$. If the group action $\phi$ is transitive, then such a lift always exists.
Theorem IV.4
Define $\Lambda_{1}:\mathcal{M}\times L\to\mathfrak{se}_{2}(3)$ as
$\Lambda_{1}(\xi,\gamma)=\gamma_{1}^{\wedge}-\boldsymbol{b}^{\wedge}+\operatorname{Ad}_{\boldsymbol{T}^{-1}}\gamma_{2}^{\wedge}+\boldsymbol{T}^{-1}f_{1}^{0}(\boldsymbol{T}),$
$\Lambda_{2}:\mathcal{M}\times L\to\mathfrak{se}_{2}(3)$ as
$\Lambda_{2}(\xi,\gamma)=\operatorname{ad}_{\boldsymbol{b}^{\wedge}}(\Lambda_{1}(\xi,\gamma))-\gamma_{3}^{\wedge},$
$\Lambda_{3}:\mathcal{M}\times L\to\mathfrak{se}(3)$ as
$\Lambda_{3}(\xi,\gamma)=\operatorname{Ad}_{\boldsymbol{K}^{-1}}\bigl{[}\Gamma(\Lambda_{1}(\xi,\gamma))\bigr{]}+\gamma_{4}^{\wedge},$
then $\Lambda:\mathcal{M}\times L\to\mathfrak{g}=\mathfrak{se}_{2}(3)\oplus\mathfrak{se}_{2}(3)\oplus\mathfrak{se}(3)$
$\Lambda(\xi,\gamma)=(\Lambda_{1}(\xi,\gamma),\Lambda_{2}(\xi,\gamma),\Lambda_{3}(\xi,\gamma))$ (8)
is an equivariant lift [19].
Let $X\in G$ be the state on the Lie group, and $\xi^{0}\in\mathcal{M}$ be the original state. If the group action $\phi$ is free, then the lift $\Lambda$ leads to a system on Lie group $G$:
$\dot{X}=dL_{X}\circ\Lambda(\phi^{(\xi^{0})}(X),\gamma).$ (9)
This transformation moves the estimation problem from the manifold to the Lie group, allowing the error to be defined via multiplication on the Lie group.
If the estimated error covariance reflects the true distribution of the errors, then the estimator is said to be consistent. The consistency of the system can be proven using the group action $\phi$ in equation (5) and the lift $\Lambda$ in equation (8) [23, 24].
V Equivariant Filter
V-A Error Dynamics
For a state $\xi\in\mathcal{M}$, define $e=\phi_{\hat{X}^{-1}}(\xi)\in\mathcal{M}$ to represent the error on the manifold $\mathcal{M}$ [25]. Choose a local coordinate map $\varphi:U\to\mathbb{R}^{9}\oplus\mathbb{R}^{9}\oplus\mathbb{R}^{6}$ at the origin $\xi^{0}$, $\varphi$ can be taken as
$\varphi(e)=(\log^{\vee}e_{1},e_{2}^{\vee},\log^{\vee}e_{3}).$
Let $\varepsilon=\varphi(e)\in\mathbb{R}^{24}$. The linear error $\varepsilon$ satisfies the differential equation
$\dot{\varepsilon}$ $\approx\boldsymbol{F}\varepsilon,$
$\boldsymbol{F}$ $=(d\varphi)_{\xi^{(0)}}\circ\bigl{(}d\phi^{(\xi^{0})}\bigr{)}_{I}\circ\bigl{(}d\Lambda_{\boldsymbol{u}}\bigr{)}_{\xi^{0}}\circ\bigl{(}d\varphi^{-1}\bigr{)}_{\boldsymbol{0}}.$
The closed-form of the matrix $\boldsymbol{F}$ is
$\boldsymbol{F}=\begin{bmatrix}\boldsymbol{F}_{T}&-\boldsymbol{I}_{9\times 9}&\boldsymbol{O}_{9\times 6}\\ \boldsymbol{O}_{9\times 9}&\boldsymbol{F}_{b}&\boldsymbol{O}_{9\times 6}\\ \boldsymbol{O}_{6\times 9}&\boldsymbol{O}_{6\times 9}&\boldsymbol{F}_{K}\end{bmatrix}\in\mathbb{R}^{24\times 24},$ (10)
where
[…]
$\boldsymbol{Y}=\bigl{(}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{a}-\hat{\boldsymbol{b}}_{a})+(\hat{\boldsymbol{v}}_{wb}^{w})^{\wedge}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\omega}-\hat{\boldsymbol{b}}_{g})+\boldsymbol{g}^{w}\bigr{)}^{\wedge},$
$\boldsymbol{Z}=\bigl{(}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\mu}-\hat{\boldsymbol{b}}_{\mu})+(\hat{\boldsymbol{r}}_{wb}^{w})^{\wedge}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\omega}-\hat{\boldsymbol{b}}_{g})+\hat{\boldsymbol{v}}_{wb}^{w}\bigr{)}^{\wedge},$
$\boldsymbol{W}=\bigl{(}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\omega}-\hat{\boldsymbol{b}}_{g})\bigr{)}^{\wedge}.$
The letters with hat in them represent the estimated value of that state and $\boldsymbol{I}$ represents the identity matrix.
V-B Measurement Model
[…]
Figure 2: The measurement model.
At a certain moment, the LiDAR acquires $m$ sampling points $\{\boldsymbol{p}_{j}\}_{j=1}^{m}$. For each point $\boldsymbol{p}_{j}$, We denote point $\boldsymbol{p}_{j}$ transformed to the world coordinate system as $\boldsymbol{p}_{j}^{w}:=\boldsymbol{C}_{b}^{w}(\boldsymbol{C}^{b}\boldsymbol{p}_{j}+\boldsymbol{l}^{b})+\boldsymbol{r}_{wb}^{w}$. Search for the five points closest to $\boldsymbol{p}_{j}^{w}$ and fit a plane, then $\boldsymbol{p}_{j}^{w}$ should be in this plane (Fig. 2). Therefore, the system output $h_{j}:\mathcal{M}\to\mathbb{R}$ can be configured as [14]
$h_{j}(\xi)=\boldsymbol{n}_{j}^{\top}\bigl{(}\boldsymbol{p}_{j}^{w}-\boldsymbol{q}_{j}\bigr{)},$
where $\boldsymbol{n}_{j}$ is the normal vector of the plane and $\boldsymbol{q}_{j}$ is an arbitrary point in the plane.
According to [8], the linearized measurement equation is
$\boldsymbol{z}_{j}=\boldsymbol{H}_{j}\varepsilon,\ \boldsymbol{H}_{j}=(dh_{j})_{\hat{\xi}}\circ(d\phi_{\hat{X}})_{\xi^{0}}\circ(d\varphi^{-1})_{\boldsymbol{0}}.$
The closed-form of the matrix $\boldsymbol{H}_{j}$ is
$\boldsymbol{H}_{j}=\begin{bmatrix}\boldsymbol{H}_{j,T}&\boldsymbol{O}_{1\times 9}&\boldsymbol{H}_{j,K}\end{bmatrix}\in\mathbb{R}^{1\times 24},$ (11)
where
$\boldsymbol{H}_{j,T}$ $=\begin{bmatrix}-\boldsymbol{n}_{j}^{\top}{\boldsymbol{p}_{j}^{w}}^{\wedge}&\boldsymbol{O}_{1\times 3}&\boldsymbol{n}_{j}^{\top}\end{bmatrix}\in\mathbb{R}^{1\times 9},$
$\boldsymbol{H}_{j,K}$ $=\begin{bmatrix}-\boldsymbol{n}_{j}^{\top}{\boldsymbol{p}_{j}^{w}}^{\wedge}&\boldsymbol{n}_{j}^{\top}\end{bmatrix}\in\mathbb{R}^{1\times 6}.$
V-C Gravity Constraints on Manifolds $\mathbb{S}^{2}$
[…]
Figure 3: Illustration of the error state on $\mathbb{S}^{2}$. The neighborhood at $\boldsymbol{g}_{k}\in\mathbb{S}^{2}$ is homeomorphic to $\mathbb{R}^{2}$. $\boldsymbol{x}$ is a minimal parameterization of the error between $\boldsymbol{g}_{k+1}$ with $\boldsymbol{g}_{k}$.
We consider gravity to be a vector with constant magnitude and direction on the sphere $\mathbb{S}^{2}$. Since $\mathbb{S}^{2}$ is a two-dimensional manifold, we use a two-dimensional vector in the tangent space $T_{\boldsymbol{g}}\mathbb{S}^{2}\simeq\mathbb{R}^{2}$ as the error state for optimization, which uses the fewest degrees of freedom to estimate gravity.
Define the direction vector of gravity as $\boldsymbol{g}=(x,y,z)^{\top}\in\mathbb{S}^{2}$ and $\left\|\boldsymbol{g}\right\|=1$. Consider the linear map [26, 27]
$B_{\boldsymbol{g}}:\mathbb{R}^{2}\to\mathbb{R}^{3}\quad B_{\boldsymbol{g}}=\begin{bmatrix}1-\frac{x^{2}}{1+z}&-\frac{xy}{1+z}\\[2.84526pt] -\frac{xy}{1+z}&1-\frac{y^{2}}{1+z}\\ -x&-y\end{bmatrix}.$ (12)
$B_{\boldsymbol{g}}$ is an isometry operator, that is, $\left\|B_{\boldsymbol{g}}\boldsymbol{x}\right\|=\left\|\boldsymbol{x}\right\|$ for all $\boldsymbol{x}\in\mathbb{R}^{2}$. Through $B_{\boldsymbol{g}}$, we can convert the vector in the tangent space $T_{\boldsymbol{g}}\mathbb{S}^{2}\simeq\mathbb{R}^{2}$ into a rotation vector, thereby updating the gravity. The pseudoinverse of the matrix $B_{\boldsymbol{g}}$ is $B_{\boldsymbol{g}}^{\top}$.
Denote the direction of gravity at time $k$ by $\boldsymbol{g}_{k}\in\mathbb{S}^{2}$ and the error state of the optimal estimate by $\boldsymbol{x}\in\mathbb{R}^{2}$. As shown in Fig. 3, we take
$\boldsymbol{g}_{k+1}=\exp(B_{\boldsymbol{g}_{k}}\boldsymbol{x})\boldsymbol{g}_{k}\in\mathbb{S}^{2}.$
Conversely, for the two gravity directions $\boldsymbol{g}_{k}$ and $\boldsymbol{g}_{k+1}$, the error is defined as
$\varepsilon_{g}=\arccos(\boldsymbol{g}_{k}\cdot\boldsymbol{g}_{k+1})B_{\boldsymbol{g}_{k}}^{\top}(\boldsymbol{g}_{k}\times\boldsymbol{g}_{k+1}),$ (13)
where $\boldsymbol{g}_{k}\cdot\boldsymbol{g}_{k+1}$ represents the dot product of the vectors.
VI Experiments
In this section, we conduct a series of experiments to evaluate the accuracy and robustness of Eq-LIO. In all experiments, we compare Eq-LIO with the IEKF-based LIO algorithm and the EKF-based FAST-LIO2 algorithm. Furthermore, for fairness, all algorithms use the same tuning parameters in the experiments.
VI-A Public Datasets
1) Public LiLi-OM Dataset: Given LiLi-OM dataset[28] has no precise ground truth, we use the start-end distance as an indicator to roughly estimate accuracy of LIO. It is worth noting that the reference value is calculated by GNSS positioning results. The distance error for the LiLi-OM dataset is presented in Table I. The results show that the proposed Eq-LIO achieves the best accuracy on the three sequences, whereas the accuracy of FAST-LIO2 and IEKF on the Schloss-2 dataset decreases significantly. In addition, Fig. 4 shows the time series of the three-axis gyroscope bias in the Schloss-1 dataset. Interestingly, the proposed Eq-LIO exhibits the ability of the filter to converge quickly.
TABLE I: Distance Errors on the LiLi-OM Dataset
{tblr}
colspec=*4X[c,m], width=hline1,Z=1pt, hline2, cell2-Z1=font= Error(m) & FAST-LIO2 IEKF EqF
Schloss-1 0.60 0.57 0.53
Schloss-2 11.50 8.75 7.07
Campus-1 0.97 0.78 0.46
Figure 4: The estimated three-axis gyroscope bias sequence in the Schloss-1 dataset.
2) Public R3LIVE Dataset: In the R3LIVE[29] dataset, we use datasets with closed trajectories so that we can evaluate the end-to-end error. The end-to-end error results we obtained are shown in Table II. In most scenarios, Eq-LIO delivers superior results. However, in certain cases, such as the sequence hkust_campus_01, the scan-to-map matching strategy used by FAST-LIO2 improves loop accuracy but also introduces the risk of matching errors, which will cause large jumps in the trajectory, as illustrated in Fig.5, while Eq-LIO does not have this problem. The speed curve is used here because it is easier to observe such jumps.
TABLE II: End-to-End Errors on the R3LIVE Dataset
{tblr}
colspec=X[1.2,c,m]*3X[c,m], width=hline1,Z=1pt, hline2, cell2-Z1=font= Error(m) & FAST-LIO2 IEKF EqF
hku_campus_00 0.17 0.07 0.06
hkust_campus_00 3.10 3.59 2.26
hkust_campus_01 0.13 1.51 0.19
hkust_campus_02 0.04 0.03 0.09
degenerate_00 8.25 6.23 6.02
Figure 5: Time series plot of the estimated velocities of the three axes in the hkust_canpus_01 dataset.
Figure 6: Estimated attitude sequences in the indoor experiment.
VI-B Private Datasets
1) Indoor Experiment: Robustness is a crucial characteristic of filter-based LiDAR-inertial odometry, referring to the system’s ability to withstand adverse external disturbances such as imperfect tuning parameters, degenerate motion scenes, or sustained intense movement. To induce significant changes in both rotation and velocity, the data collectors held the sensor in their hands and ran irregularly indoors, generating large and unpredictable movements. The data were collected using a Livox Avia LiDAR and its built-in MEMS IMU, with sampling frequencies of 100 Hz for LiDAR and 10 Hz for the IMU. Fig. 6 illustrates the attitude estimation during the experiment, revealing dramatic changes in a short period, making it ideal for examining the robustness of the proposed system. The results, depicted in Fig. 7, demonstrate that Eq-LIO produces the clearest mapping result and handles wall corners more effectively than IEKF. The mapping results of EKF-based FAST-LIO2 exhibit significant drift, while Eq-LIO shows less distortion on the walls, with the upper right corner of the map appearing clearer. This suggests that when handling highly nonlinear motion scenarios, the model based on Eq-LIO exhibits a superior linearization effect, making it more resilient to significant parameter errors and demonstrating enhanced robustness.
(a) Eq-LIO
(b) IEKF
(c) FAST-LIO2
Figure 7: Mapping results of LIO in the indoor high-speed environment based on three different filtering algorithms. The mapping results of FAST-LIO2 for this data set are not available. The red lines depict the wall of the EqF and IEKF mapping results, which can reflect the degree of deformation of the wall.
2) Outdoor Experiment: To verify the performance of Eq-LIO in actual urban complex scenes, we collected data from real urban highway scenes. The dataset was collected on roads in Wuhan and includes challenging scenarios such as long straight lines and large arc U-turns. The presence of many moving vehicles further complicates achieving robust navigation. The hardware platform for data collection is as follows: a GNSS receiver (Septentrio PolaRx5), GNSS antenna (Trimble Zephyr Model2), a MEMS IMU (ADIS-16470), a tactical IMU (StarNeto XW-GI7660), and a mechanical LiDAR (Velodyne VLP-32) were rigidly fixed on an iron plate with roughly pre-calibrated extrinsic parameters. We used the MEMS IMU for testing and the tactical IMU with RTK/INS to generate reference ground truth.
Figure 8: Trajectories estimated by three filter-based LIO algorithms in the outdoor experiment.
TABLE III: Planar Distance Errors at the End of the Trajectory in the Outdoor Experiment.
{tblr}
colspec=X[0.8,c,m]*3X[c,m], width=hline1,Z=1pt, hline2 & E(m) N(m) Plane(m)
FAST-LIO2 14.309 51.725 53.668
IEKF 25.079 60.302 65.309
Eq-LIO 13.899 25.124 28.713
As shown in Fig. 8, all algorithms exhibit some errors due to the long trajectory and the presence of numerous moving objects. However, at both the U-turn and the end of the trajectory, the proposed Eq-LIO demonstrates superior performance, accurately estimating the scale of trajectory. Table III shows the planar distance errors of different algorithms at the final moment. The trajectory length is approximately 5132 meters, with Eq-LIO exhibiting the smallest drift in both the east and north directions. The vehicle primarily traveled along the north-south axis, resulting in more significant inconsistency in the northward direction. The results indicate that Eq-LIO produces smaller errors under erroneous observations compared to IEKF and EKF, demonstrating superior consistency in the northward direction.
VII Conclusion
This article presents Eq-LIO. To address the inconsistency and linearization point issues of traditional filtering algorithms, this work develops an equivariant filter for LIO systems, derives a new error definition based on the symmetry of the semi-direct product group, and incorporates gravity constraints. The algorithm achieves improved accuracy and robustness without increasing computational resources. Experiments on various public and private datasets show that, compared to existing methods, Eq-LIO can achieve higher accuracy and more importantly, robustness to challenging motion scenes at comparable speeds.